# Action-Driven Flows for Causal Variational Principles

### v1

@@ 4.2. Minimizing movements
for $j=1,2,\ldots$. The necessity of introducing an additional penalization term depending on $\xi$ will be addressed in Section 4.4.
@@ 4.3. A Hölder continuous flow
The curve $\rho^{h,\xi}(t)$ defined by (4.9) satisfies for all $0<t_{1}<t_{2}<\infty$ the inequality
$d\big{(}\rho^{h,\xi}(t_{1}),\rho^{h,\xi}(t_{2})\big{)}\leqslant\sqrt{2}\>\sqrt{t_{2}-t_{1}+h}\;\sqrt{{\mathcal{S}}(\rho_{0})}\>.$ (4.12)
@@
Let $0<t_{1}<t_{2}<\infty$. Then, by definition of $\rho_{h,\xi}$,
@@
$\qquad\qquad+\sum_{j=\lfloor\frac{t_{1}}{h}+1\rfloor}^{\lfloor\frac{t_{2}}{h}\rfloor-1}d\big{(}\rho_{j}^{h,\xi},\rho_{j+1}^{h,\xi}\big{)}$
@@
$\leqslant\bigg{(}\sum_{j=\lfloor\frac{t_{1}}{h}\rfloor}^{\lfloor\frac{t_{2}}{h}+1\rfloor-1}1\bigg{)}^{\frac{1}{2}}\bigg{(}\sum_{j=\lfloor\frac{t_{1}}{h}\rfloor}^{\lfloor\frac{t_{2}}{h}+1\rfloor-1}2h\>\big{(}{\mathcal{S}}(\rho_{j}^{h,\xi})-{\mathcal{S}}(\rho_{j+1}^{h,\xi})\big{)}\bigg{)}^{\frac{1}{2}}$
(4.6)${\leqslant}\sqrt{2}\>\sqrt{(t_{2}-t_{1})+h}\>\sqrt{{\mathcal{S}}(\rho_{0})}\>,$
where the second step is trivial for $d$ being the Fréchet metric and follows from Lemma 2.1 in the case of the Wasserstein metric. This completes the proof. ∎
We are now ready for proving our first main result.
Theorem 4.6.
@@
Case 1. $d=W_{p}$. Let $[T_{1},T_{2}]\subset[0,\infty)$ be a compact interval. We note that $(\mathfrak{M}_{1}(\mathscr{F}),W_{p})$ is a compact, hence complete, metric space by the Banach-Alaoglu theorem. We aim to apply Lemma 4.4 to the sequence $(\rho^{\xi,1/j})_{j\in\mathbb{N}}$ together with $d=W_{p}$ and $\tau$ being the weak*-topology on $\mathfrak{M}_{1}(\mathscr{F})$. Then $\rho^{\xi,1/j}(t)\in K:=\mathfrak{M}_{1}(\mathscr{F})$ for all $j\in\mathbb{N}$, whereby (4.10) is satisfied. Moreover, the estimate (4.12) yields that (4.11) is fulfilled with $\omega(s,t):=\sqrt{2|s-t|}$. Consequently, Lemma 4.4 gives the existence of a $W_{p}$-continuous limit map $\rho^{\xi}\colon[T_{1},T_{2}]\to\mathfrak{M}_{1}(\mathscr{F})$ such that $\rho^{\xi,1/j(k)}\to\varrho^{\xi}$ pointwise everywhere in $[T_{1},T_{2}]$ with respect to $d=W_{p}$. For all $T_{1}\leqslant t_{1}\leqslant t_{2}\leqslant T_{2}$ we thus obtain
@@
In order to construct the requisite curve as claimed in Theorem 4.6, we cover $[0,\infty)$ by intervals $I_{\ell}:=[\ell-1,\ell+1]$, $\ell\in\mathbb{N}$. By what has been said above, we may choose a sequence $(j_{k}^{(1)})$ such that, for a certain limit curve $\varrho^{\xi}\in\operatorname{C}^{0,1/2}([0,2];\mathfrak{M}_{1}(\mathscr{F}))$ we have
@@
Case 2. $d=d_{\mathfrak{M}(\mathscr{F})}$. In this situation, we let $d=d_{\mathfrak{M}(\mathscr{F})}$ and again let $\tau$ be the weak*-topology on $\mathfrak{M}_{1}(\mathscr{F})$. Then $K:=\mathfrak{M}_{1}(\mathscr{F})$ is compact for $\tau$. Arguing as above, specifically applying (4.12) to $d=d_{\mathfrak{M}(\mathscr{F})}$, we obtain the existence of a limit map $\varrho^{\xi}\in[0,\infty)\in\operatorname{C}([0,\infty);(\mathfrak{M}_{1}(\mathscr{F});d_{\mathfrak{M}(\mathscr{F})}))$ such that, for some sequence $(j_{l})$ with $j_{l}\to\infty$ as $l\to\infty$, $\rho^{\xi,1/j_{l}}\to\varrho$ uniformly in $(\mathfrak{M}_{1}(\mathscr{F}),W_{p})$ (not in $(\mathfrak{M}_{1}(\mathscr{F}),d_{\mathfrak{M}(\mathscr{F})})$) on compact subsets of $[0,\infty)$.
Let us note that we have $\varrho^{\xi}\in\operatorname{C}_{\operatorname{loc}}^{0,1/2}([0,\infty);(\mathfrak{M}_{1}(\mathscr{F}),d_{\mathfrak{M}(\mathscr{F})}))$ indeed: Let $0\leqslant T_{1}\leqslant T_{2}<\infty$, so that $\rho^{\xi,1/j_{l}}(t)\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\varrho^{\xi}(t)$ for all $t\in[T_{1},T_{2}]$ since $W_{p}$ metrizes weak*-convergence on $\mathfrak{M}(\mathscr{F})$. Since in the present setting (4.12) is available for $d=d_{\mathfrak{M}(\mathscr{F})}$, we conclude for $t_{1},t_{2}\in[T_{1},T_{2}]$ by weak*-lower semicontinuity of the total variation norm
@@
In this sense, the passage to the weak*-metric is only required to obtain the existence of such a curve, whereas the Hölder regularity for $d_{\mathfrak{M}(\mathscr{F})}$ survives from Lemma 4.5 by lower semicontinuity. This concludes the proof of Theorem 4.6. ∎
@@ 4.4. A Lipschitz curve in the case $\xi>0$
showing that the length of the discrete curve is bounded by the total change of the action. This estimate suggests that it is useful to use the action itself for the parametrization of the curve. As we shall see, it is of advantage to do so already for the discrete curve, before taking the limit $h\searrow 0$ (as will be explained Remark 4.13 below). To this end, given $h,\xi>0$, we set
@@
$\tilde{\rho}^{h,\xi}(s):=\frac{s_{j}-s}{s_{j}-s_{j+1}}\>\rho^{h,\xi}_{j}+\frac{s_{j+1}-s}{s_{j}-s_{j+1}}\>\rho^{h,\xi}_{j+1}\qquad\text{if~{}$s\in[s_{j+1},s_{j}]$}\>.$ (4.16)
@@
Assume that the Lagrangian is Hölder continuous, ${\mathcal{L}}\in\operatorname{C}^{0,\alpha}(\mathscr{F}\times\mathscr{F},\mathbb{R}^{+}_{0})$. Then there is a constant $C>0$ (which depends only on $\mathscr{F}$ and ${\mathcal{L}}$) such that for all $s,s^{\prime}\in\big{[}{\mathcal{S}}\big{(}\rho^{h,\xi}_{\infty}\big{)},{\mathcal{S}}\big{(}\rho_{0}\big{)}\big{]}$ and $h>0$,
@@
Then the triangle inequality yields
@@
$\frac{1}{\xi}\>\big{|}s_{j}-s\big{|}=\frac{1}{\xi}\>\Big{|}{\mathcal{S}}\big{(}\tilde{\rho}^{h,\xi}(s)\big{)}-{\mathcal{S}}\big{(}\tilde{\rho}^{h,\xi}(s)\big{)}\Big{|}\leqslant\frac{C}{\xi}\>d\big{(}\tilde{\rho}^{h,\xi}(s),\tilde{\rho}^{h,\xi}(s)\big{)}^{\alpha}\>,$
and again applying 2.1 and (4.4) gives
@@
By iteratively choosing subsequences and taking the limit of the diagonal sequences, one obtains a curve of measures denoted by
@@
${\mathcal{S}}^{\xi}_{\min}:=\liminf_{h\searrow 0}{\mathcal{S}}\big{(}\rho^{h,\xi}_{\infty}\big{)}\>.$
@@
$\tilde{\rho}^{h_{n_{k}},\xi}(s_{\ell})\overset{k\rightarrow\infty}{\longrightarrow}\tilde{\varrho}^{\xi}(s_{\ell})\qquad\text{for all~{}$\ell\in\mathbb{N}$}\>.$
Considering the estimate (4.7) and passing to the limit, we find that the family of limit measures is again Lipschitz continuous in the sense that
@@
In order to prove (4.19), we estimate the Wasserstein distance (which, as specified in (2.4), metrizes the weak*-topology). We first note that, for any $\ell\in\mathbb{N}$,
@@
$\liminf_{k\rightarrow\infty}W_{p}\Big{(}\tilde{\rho}^{h_{n_{k}},\xi}\big{(}{\mathcal{S}}\big{(}\rho^{h,\xi}_{\infty}\big{)}\big{)},\tilde{\varrho}^{\xi}\big{(}{\mathcal{S}}^{\xi}_{\min}\big{)}\Big{)}\leqslant\frac{2}{\xi}\>\Big{(}{\mathcal{S}}^{\xi}_{\min}-s_{\ell}\Big{)}\>.$
@@ 4.5. Limiting measures and Euler-Lagrange equations
In the case $\xi=0$, assume that the curve $\varrho^{0}(t)$ with initial measure $\varrho^{0}(0)=\rho_{0}\in\mathfrak{M}_{1}(\mathscr{F})$ converges in the weak*-sense. We set
@@
In the case $\xi>0$, the curve $\varrho^{\xi}(s)$ converges as $s\searrow{\mathcal{S}}^{\xi}_{\min}$. In the case of penalization by the Wasserstein distance $W_{p}$ (i.e., Case 2 in (4.2)), the limiting measure
@@
is minimal on the support of $\varrho^{\xi}_{\infty}$,
@@
so that
@@
×$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$ $\leqslant\int_{\mathscr{F}}{\mathcal{L}}(x,z)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)+\frac{\xi}{2}\>W_{p}(\delta_{z},\rho_{\infty}^{h,\xi})\>.$ (4.21)
@@
Let $10<\tau<$ 1, $z\in\mathscr{F}$ and define
$\mu_{\tau}^{j,h,\xi}:=(1-\tau)\rho_{j}^{h,\xi}+\tau\delta_{z}\in\mathfrak{M}_{1}(\mathscr{F}).$
Minimality consequently implies that
00 0 $\leqslant\frac{1}{\tau}\Big{(}\mathcal{S}(\mu_{\tau}^{j,h,\xi})-\mathcal{S}(\rho_{j}^{h,\xi})\Big{)}+\frac{1}{2\tau h}\Big{(}W_{p}(\mu_{\tau}^{j,h,\xi},\rho_{j-1}^{h,\xi})^{2}-W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})^{2}\Big{)}$
@@
.$\Big{.}\;\;\;\;\;\;\;\;+2(1-\tau)\>W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})\>W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})+\tau W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})^{2}\Big{)}\>.$
Since $\xi>0$ and $h>0$ are fixed, we have that $W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})\to 0$ as $j\to\infty$. Moreover, $\sup_{j\in\mathbb{N}}W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})<\infty$, and therefore
@@
00 0 $\leqslant\frac{(1-\tau)^{2}}{\tau}\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)-\frac{1}{\tau}\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$
@@
×$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$ $\leqslant\int_{\mathscr{F}}{\mathcal{L}}(x,z)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)+\frac{\xi}{2}W_{p}(\delta_{z},\rho_{\infty}^{h,\xi})$
$=\inf_{\mathscr{F}}\ell_{\infty}^{h,\xi}+\frac{\xi}{2}W_{p}(\delta_{z},\rho_{\infty}^{h,\xi}).$
@@
×$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\mu(x)\operatorname{d}\!\mu(y)\leqslant\inf_{y\in\mathscr{F}}\int_{\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\mu(x)+\frac{\xi}{2}\>W_{p}(\delta_{z},\mu)\>.$ (4.26)
@@
According to Lemma 4.12, it suffices to show that there is a sequence $(h_{k})_{k\in\mathbb{N}}$ with $h_{k}\searrow 0$ such that the corresponding discrete limit measures $(\rho_{\infty}^{h_{k},\xi})$ and converge to the limit measure $\varrho_{\infty}$ respectively $\varrho^{\xi}_{\infty}$. In the case $\xi=0$, this is a consequence of the assumptions in Theorem 4.9. In the case $\xi>0$, on the other hand, this was proved in (4.19). ∎

### v2

@@ 4.2. Minimizing movements
for $j=1,2,\ldots$. The first penalization term follows the general procedure in the minimizing movements approach (see for example [2]); also the resulting Hölder estimates (as in Lemma 4.5 and Proposition 4.6) are adaptations of standard arguments to our setting (see for example [4, Proposition 7.1]). The second penalization term in (4.5), however, is novel. The necessity of introducing this additional penalization term depending on $\xi$ will be explained in detail in Section 4.4.
@@ 4.3. A Hölder continuous flow
The curve $\rho^{h,\xi}(t)$ defined by (4.9) satisfies for all $0<t_{1},t_{2}<\infty$ the inequality
$d\big(\rho^{h,\xi}(t_{1}),\rho^{h,\xi}(t_{2})\big)\leqslant\sqrt{2}\>\sqrt{|t_{2}-t_{1}|+h}\;\sqrt{{\mathcal{S}}(\rho_{0})}\>.$ (4.12)
@@
It clearly suffices to consider the case $t_{1}<t_{2}$. Then, by definition of $\rho_{h,\xi}$,
@@
$\qquad\qquad+\sum_{j=\lfloor\frac{t_{1}}{h}+1\rfloor}^{\lfloor\frac{t_{2}}{h}\rfloor-1}d\big(\rho_{j}^{h,\xi},\rho_{j+1}^{h,\xi}\big)\>,$
where the last step is trivial for $d$ being the Fréchet metric and follows from Lemma 2.1 in the case of the Wasserstein metric. It follows that
$d\big(\rho^{h,\xi}(t_{1}),\rho^{h,\xi}(t_{2})\big)\leqslant\sum_{j=\lfloor\frac{t_{1}}{h}\rfloor}^{\lfloor\frac{t_{2}}{h}+1\rfloor-1}d(\rho_{j}^{h,\xi},\rho_{j+1}^{h,\xi})$
@@
$\leqslant\bigg(\sum_{j=\lfloor\frac{t_{1}}{h}\rfloor}^{\lfloor\frac{t_{2}}{h}+1\rfloor-1}1\bigg)^{\frac{1}{2}}\bigg(\sum_{j=\lfloor\frac{t_{1}}{h}\rfloor}^{\lfloor\frac{t_{2}}{h}+1\rfloor-1}2h\>\big({\mathcal{S}}(\rho_{j}^{h,\xi})-{\mathcal{S}}(\rho_{j+1}^{h,\xi})\big)\bigg)^{\frac{1}{2}}\>.$
The last sum is telescopic. Moreover, using that the sequence of actions is monotone decreasing (4.6) and non-negative, we conclude that
$d\big(\rho^{h,\xi}(t_{1}),\rho^{h,\xi}(t_{2})\big)\leqslant\sqrt{2}\>\sqrt{(t_{2}-t_{1})+h}\>\sqrt{{\mathcal{S}}(\rho_{0})}\>.$
This completes the proof. ∎
We are now ready for proving our first existence result.
Proposition 4.6.
@@
Case 1. $d=W_{p}$. Let $T_{1},T_{2}]\subset[0,\infty)$ be a compact interval. We note that $(\mathfrak{M}_{1}(\mathscr{F}),W_{p})$ is a compact, hence complete, metric space by the Banach-Alaoglu theorem. We aim to apply Lemma 4.4 to the sequence $(\rho^{\xi,1/j})_{j\in\mathbb{N}}$ together with $d=W_{p}$ and $\tau$ being the weak*-topology on $\mathfrak{M}_{1}(\mathscr{F})$. Then $\rho^{\xi,1/j}(t)\in K:=\mathfrak{M}_{1}(\mathscr{F})$ for all $j\in\mathbb{N}$, whereby (4.10) is satisfied. Moreover, the estimate (4.12) yields that (4.11) is fulfilled with $\omega(s,t):=\sqrt{2|s-t|}$. Consequently, Lemma 4.4 together with (2.5) gives the existence of a $W_{p}$-continuous limit map $\rho^{\xi}\colon[T_{1},T_{2}]\to\mathfrak{M}_{1}(\mathscr{F})$ such that $\rho^{\xi,1/j(k)}(t)\to\varrho^{\xi}(t)$ with respect to $d=W_{p}$ for every $t\in[T_{1},T_{2}]$. For all $T_{1}\leqslant t_{1}\leqslant t_{2}\leqslant T_{2}$ we thus obtain
@@
In order to construct the requisite curve as claimed in Proposition 4.6, we cover $0,\infty)$ by intervals $I_{\ell}:=[\ell-1,\ell+1]$, $\ell\in\mathbb{N}$. By what has been said above, we may choose a sequence $(j_{k}^{(1)})$ such that, for a certain limit curve $\varrho^{\xi}\in\operatorname{C}^{0,1/2}([0,2];\mathfrak{M}_{1}(\mathscr{F}))$ we have
@@
Case 2. $d=d_{\mathfrak{M}(\mathscr{F})}$. In this situation, we let $d=d_{\mathfrak{M}(\mathscr{F})}$ and again let $\tau$ be the weak*-topology on $\mathfrak{M}_{1}(\mathscr{F})$. Then $K:=\mathfrak{M}_{1}(\mathscr{F})$ is compact for $\tau$. Arguing as above, specifically applying (4.12) to $d=d_{\mathfrak{M}(\mathscr{F})}$, we obtain the existence of a limit map $\varrho^{\xi}\in\operatorname{C}([0,\infty);(\mathfrak{M}_{1}(\mathscr{F});d_{\mathfrak{M}(\mathscr{F})}))$ such that, for some sequence $(j_{l})$ with $j_{l}\to\infty$ as $l\to\infty$, $\rho^{\xi,1/j_{l}}\to\varrho$ in $(\mathfrak{M}_{1}(\mathscr{F}),W_{p})$ (not in $(\mathfrak{M}_{1}(\mathscr{F}),d_{\mathfrak{M}(\mathscr{F})})$), locally uniformly in time (i.e. uniformly in $t$ in a compact subset of $0,\infty)$).
Let us note that we have $\varrho^{\xi}\in\operatorname{C}_{\operatorname{loc}}^{0,1/2}([0,\infty);(\mathfrak{M}_{1}(\mathscr{F}),d_{\mathfrak{M}(\mathscr{F})}))$ indeed: Let $0\leqslant T_{1}\leqslant T_{2}<\infty$, so that $\rho^{\xi,1/j_{l}}(t)\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\varrho^{\xi}(t)$ for all $t\in[T_{1},T_{2}]$ since $W_{p}$ metrizes weak*-convergence on $\mathfrak{M}(\mathscr{F})$. Since in the present setting (4.12) is available for $d=d_{\mathfrak{M}(\mathscr{F})}$, we conclude for $t,t^{\prime}\in[T_{1},T_{2}]$ by weak*-lower semicontinuity of the total variation norm
@@
In this sense, the passage to the weak*-metric is only required to obtain the existence of such a curve, whereas the Hölder regularity for $d_{\mathfrak{M}(\mathscr{F})}$ survives from Lemma 4.5 by lower semicontinuity. This concludes the proof of Proposition 4.6. ∎
@@ 4.4. A Lipschitz curve in the case $\xi>0$
showing that the length of the discrete curve is bounded by the total change of the action. This estimate suggests that it is useful to use the action itself for the parametrization of the curve. As we shall see, it is of advantage to do so already for the discrete curve, before taking the limit $h\searrow 0$ (as will be explained in Remark 4.13 below). To this end, given $h,\xi>0$, we set
@@
$\tilde{\rho}^{h,\xi}(s):=\frac{s_{j}-s}{s_{j}-s_{j+1}}\>\rho^{h,\xi}_{j}+\frac{s-s_{j+1}}{s_{j}-s_{j+1}}\>\rho^{h,\xi}_{j+1}\qquad\text{if~$s\in[s_{j+1},s_{j}]$}\>.$ (4.16)
@@
[…]
@@
Applying the triangle inequality as well as (4.7) yields
@@
$\frac{1}{\xi}\>\big|s_{j}-s\big|=\frac{1}{\xi}\>\Big|{\mathcal{S}}\big(\tilde{\rho}^{h,\xi}(s_{j})\big)-{\mathcal{S}}\big(\tilde{\rho}^{h,\xi}(s)\big)\Big|\leqslant\frac{C}{\xi}\>d\big(\tilde{\rho}^{h,\xi}(s_{j}),\tilde{\rho}^{h,\xi}(s)\big)^{\alpha}\>.$
Again Applying Lemma 2.1 and (4.4) gives
@@
By iteratively choosing subsequences and taking the limit of the diagonal sequence, one obtains a curve of measures denoted by
@@
${\mathcal{S}}^{\xi}_{\min}:=\liminf_{h\searrow 0}{\mathcal{S}}\big(\tilde{\rho}^{h,\xi}_{\infty}\big)\>.$
@@
[…]
Considering the interpolation (4.9), applying the estimate (4.7) and passing to the limit, we find that the family of limit measures is again Lipschitz continuous in the sense that
@@
In order to prove (4.19), we estimate the Wasserstein distance (which, as specified in (2.5), metrizes the weak*-topology). We first note that, for any $\ell\in\mathbb{N}$ and $h>0$,
@@
$\liminf_{k\rightarrow\infty}W_{p}\Big(\tilde{\rho}^{h_{n_{k}},\xi}\big({\mathcal{S}}\big(\rho^{h_{n_{k}},\xi}_{\infty}\big)\big),\tilde{\varrho}^{\xi}\big({\mathcal{S}}^{\xi}_{\min}\big)\Big)\leqslant\frac{2}{\xi}\>\Big({\mathcal{S}}^{\xi}_{\min}-s_{\ell}\Big)\>.$
@@ 4.5. Limiting measures and Euler-Lagrange equations
Consider the minimizing movement flow corresponding to the action with penalization (4.1) and (2.1), where the Lagrangian $\mathcal{L}$} has the properties (A1) and (A2) stated in the preliminaries on page 2.1. In the case $\xi=0$, assume that the curve $\varrho^{0}(t)$ with initial measure $\varrho^{0}(0)=\rho_{0}\in\mathfrak{M}_{1}(\mathscr{F})$ converges in the weak*-sense. We set
@@
Consider the minimizing movement flow corresponding to the action with penalization (4.1) and (2.1), where the Lagrangian $\mathcal{L}$} has the properties
(A1) and (A2) stated in the preliminaries on page 2.1. In the case $\xi>0$, the curve $\varrho^{\xi}(s)$ converges as $s\searrow{\mathcal{S}}^{\xi}_{\min}$. In the case of penalization by the Wasserstein distance $W_{p}$ (i.e., Case 2 in (4.2)), the limiting measure
@@
is minimal on $N:=\operatorname{supp}\varrho^{\xi}_{\infty}$,
@@
so that the interpolated measure defined in (4.9) can be written as
@@
$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$ $\leqslant\int_{\mathscr{F}}{\mathcal{L}}(x,z)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)+\frac{\xi}{2}\>W_{p}(\delta_{z},\rho_{\infty}^{h,\xi})\>.$ (4.21)
@@
Given $0<\tau<1$ and $z\in\mathscr{F}$ we define
$\mu_{\tau}^{j,h,\xi}:=(1-\tau)\rho_{j}^{h,\xi}+\tau\delta_{z}\in\mathfrak{M}_{1}(\mathscr{F})\>.$
Using that $\rho_{j}^{h,\xi}$ is a minimizer of the penalized action, it follows that
$0$ $\leqslant\frac{1}{\tau}\Big(\mathcal{S}(\mu_{\tau}^{j,h,\xi})-\mathcal{S}(\rho_{j}^{h,\xi})\Big)+\frac{1}{2\tau h}\Big(W_{p}(\mu_{\tau}^{j,h,\xi},\rho_{j-1}^{h,\xi})^{2}-W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})^{2}\Big)$
@@
$\Big.\;\;\;\;\;\;\;\;+2(1-\tau)\>W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})\>W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})+\tau W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})^{2}\Big)\>.$
Since $\xi\geq 0$ and $h>0$ are fixed, we have that $W_{p}(\rho_{j}^{h,\xi},\rho_{j-1}^{h,\xi})\to 0$ as $j\to\infty$. Moreover, $\sup_{j\in\mathbb{N}}W_{p}(\delta_{z},\rho_{j-1}^{h,\xi})<\infty$, and therefore
@@
$0$ $\leqslant\frac{(1-\tau)^{2}}{\tau}\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)-\frac{1}{\tau}\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$
@@
$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)\operatorname{d}\!\rho_{\infty}^{h,\xi}(y)$ $\leqslant\int_{\mathscr{F}}{\mathcal{L}}(x,z)\operatorname{d}\!\rho_{\infty}^{h,\xi}(x)+\frac{\xi}{2}W_{p}(\delta_{z},\rho_{\infty}^{h,\xi})\>.$
@@
$\iint_{\mathscr{F}\times\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\mu(x)\operatorname{d}\!\mu(y)\leqslant\inf_{y\in\mathscr{F}}\int_{\mathscr{F}}{\mathcal{L}}(x,y)\operatorname{d}\!\mu(x)+\frac{\xi}{2}\>W_{p}(\delta_{z},\mu)\>.$ (4.26)
@@
According to Lemma 4.12, it suffices to show that there is a sequence $(h_{k})_{k\in\mathbb{N}}$ with $h_{k}\searrow 0$ such that the corresponding discrete limit measures $(\rho_{\infty}^{h_{k},\xi})$ converge to the limit measure $\varrho_{\infty}$ respectively $\varrho^{\xi}_{\infty}$. In the case $\xi=0$, this is a consequence of the assumptions in Theorem 4.9. In the case $\xi>0$, on the other hand, this was proved in (4.19). ∎